# Quantitative Attractor Analysis of High-Capacity Kernel Logistic Regression Hopfield Networks

### v3

@@ Abstract
Kernel-based learning methods such as Kernel Logistic Regression (KLR) can dramatically increase the storage capacity of Hopfield networks, but the principles governing their performance and stability remain largely uncharacterized. This paper presents a comprehensive quantitative analysis of the attractor landscape in KLR-trained networks to establish a solid foundation for their design and application. Through extensive, statistically validated simulations, we address critical questions of generality, scalability, and robustness. Our comparative analysis reveals that KLR and Kernel Ridge Regression (KRR) exhibit similarly high storage capacities and clean attractor landscapes under typical operating conditions, suggesting this is a general property of kernel regression methods, though KRR is computationally much faster. We uncover a non-trivial, scale-dependent scaling law for the kernel width ($\gamma$), demonstrating that optimal capacity requires $\gamma$ to be scaled such that $\gamma\times N$ increases with network size $N$. This implies that larger networks necessitate more localized kernels—where each pattern’s influence is more spatially confined—to manage inter-pattern interference. Under this optimized scaling, we provide definitive evidence that the storage capacity scales linearly with network size ($P\propto N$). Furthermore, our sensitivity analysis shows that performance is remarkably robust to the choice of the regularization parameter $\lambda$. Collectively, these findings provide a clear set of empirical principles for designing high-capacity, robust associative memories and clarify the mechanisms that enable kernel methods to overcome the classical limitations of Hopfield-type models.
@@ 1 Introduction
Associative memory (AM) is a fundamental cognitive function that enables the retrieval of stored information based on partial or noisy cues. Artificial neural networks—particularly the Hopfield network [1]—have long served as foundational models for AM due to their biologically plausible recurrent dynamics and the interpretation of stored memories as stable states (attractors) of an energy function. The recall process in a Hopfield network can thus be viewed as the network state evolving toward a local minimum of its energy landscape, corresponding to the stored pattern most similar to the input cue.
Despite their conceptual elegance and biological grounding, traditional Hopfield networks trained with the Hebbian learning rule [1] suffer from severe limitations. Their storage capacity is theoretically bounded at a pattern-to-neuron ratio $(P/N)$ of approximately 0.14 [2]. Exceeding this limit leads to catastrophic interference, recall errors, and the proliferation of spurious attractors—states that are not correspond to any learned pattern but still act as fixed points of the network dynamics [3]. These spurious attractors compete with desired memories, shrinking their basins of attraction and severely impairing noise robustness.
To overcome these limitations, numerous approaches have reframed associative memory learning as a supervised optimization problem, where synaptic weights are tuned to make desired patterns stable fixed points. Linear approaches, such as the pseudoinverse rule [4] and Linear Logistic Regression (LLR) [5], offer moderate improvements in storage capacity and robustness. However, their performance remains fundamentally constrained by the requirement of linear separability—either in the input space or in the neuron’s input–output mapping.
Nonlinear kernel methods offer a powerful means to transcend these constraints. Techniques such as Support Vector Machines, Kernel Ridge Regression (KRR), and Kernel Logistic Regression (KLR) implicitly map data into high- or even infinite-dimensional feature spaces, where complex nonlinear relationships become linearly separable [6]. Applying this framework to associative memory learning holds the potential to dramatically increase storage capacity and enhance noise robustness.
@@ 2 Related Work
While these important lines of research—from classical KAM to MHNs—have established the potential for high-capacity associative memory, their focus has primarily been on the static properties of the energy function or performance on specific pattern reconstruction tasks. A detailed, quantitative understanding of the attractor landscape and recall dynamics that emerge from practical, supervised kernel methods like KLR and KRR remains largely unexplored. Key questions persist: What happens to spurious attractors in these high-capacity regimes? What are the primary failure modes of recall? And what are the geometric properties of the attractors themselves?
The present study aims to fill this gap. We bridge the gap between the high-level capacity analysis of KAM/MHN and the classical dynamical systems approach to Hopfield networks. By performing a detailed quantitative analysis of the attractor landscape formed by KLR—and comparing it with KRR, for which we have separately established its high performance and computational advantages [8]—we provide a microscopic, dynamical picture that explains how these kernel-based methods achieve their remarkable performance, thus complementing the existing body of work.
@@ 3.2.1 Hebbian Learning
$s_{i}(t+1)=\text{sign}\left(\sum_{j\neq i}W_{ij}s_{j}(t)\right).$ (2)
@@ 3.2.4 Kernel Ridge Regression (KRR) Hopfield Network
where $\bm{K}$ is the $P\times P$ kernel (Gram) matrix $K_{\mu\nu}=K(\bm{\xi}^{\mu},\bm{\xi}^{\nu})$, $I$ is the identity matrix, and $\bm{Y}$ is the $P\times N$ matrix of stored patterns. This non-iterative solution makes the training process of KRR significantly faster than the gradient-based optimization required for KLR, a difference we quantify in our experiments.
@@ 3.4 Recall Process
The recall process in KLR-trained networks differs from that in Hebbian and LLR-based networks, as it does not rely on an explicit $N\times N$ weight matrix. Given a network state $\bm{s}(t)$ at time $t$, the next state $\bm{s}(t+1)$ is determined by the sign of each neuron’s activation potential relative to a threshold $\theta$:
@@ 4.1 Performance and Attractor Landscape Characteristics ($N=500$)
Finally, the recall dynamics are exceptionally fast, as illustrated in Fig. 1(d). For high-similarity initial states, the network consistently converges in just one to two steps. Even for highly corrupted states in the low-similarity regime, the average number of steps to convergence remains remarkably low, typically well under 10 steps. Interestingly, for these noisy inputs, the convergence time exhibits a non-monotonic “hump” shape, peaking at intermediate storage loads. The subsequent decrease in convergence time at very high loads corresponds precisely to the regime where recall failures dominate (cf. Fig. 1(b)). This suggests that the dynamics quickly settle into nearby stable states—either incorrect learned patterns or spurious fixed points—rather than undertaking a prolonged search for the correct target. This rapid convergence is a significant practical advantage of the model.
@@ 4.2 Comparison with Kernel Ridge Regression ($N=100$)
The results of this comparison are summarized in Fig. 2. We first evaluated the recall performance under a moderate noise condition (initial similarity = 0.6), as shown in Fig. 2(a). The performance of the two methods is strikingly similar. Both KLR and KRR maintain near-perfect recall for storage loads up to $P/N\approx 3.0$. Beyond this critical capacity, both methods exhibit a sharp but graceful degradation in performance, with their recall curves closely tracking each other. While minor quantitative differences exist—with KLR showing a slightly wider shoulder in this instance—the overall qualitative behavior is virtually identical. This result strongly suggests that the high-capacity and robust attractor dynamics observed are not specific to KLR, but are likely a more general property of Hopfield-type networks trained with kernel regression methods that effectively separate patterns in a high-dimensional feature space.
@@ 5 Discussion
We hypothesize that two aspects of KLR learning are crucial. First, the kernel trick implicitly maps the input patterns into a high-dimensional feature space where they become more easily separable [6]. This enhanced separability allows the learning algorithm to find decision boundaries that create wide, well-defined basins of attraction for the stored patterns, effectively pushing potential spurious minima away from the recall trajectories. Second, the logistic loss function, combined with L2-regularization, encourages the formation of a “smooth” energy landscape. Unlike Hebbian learning, which is based on simple pairwise correlations and can create many noisy local minima, the KLR objective seeks a solution with a large margin, which tends to flatten the landscape in regions far from the decision boundaries. This landscape smoothing further inhibits the formation of unintended spurious attractors. A detailed theoretical investigation of this mechanism is the subject of our companion paper [12].
@@
Generality of High-Capacity Kernel-Based Memories: Our direct comparison between KLR and KRR (Fig. 2) yields a crucial insight: the remarkable ability to form a clean, high-capacity attractor landscape is not a feature unique to the logistic loss function of KLR. The qualitatively similar performance of KRR suggests that this is a more general property of kernel regression methods applied to Hopfield-type networks. The key mechanism appears to be the kernel trick itself, which maps patterns into a high-dimensional feature space where they become more easily separable, allowing the learning algorithm to sculpt well-defined attractor basins. While their performance is similar, the significant computational advantage of KRR’s non-iterative solution makes it a compelling practical alternative, highlighting a clear trade-off between algorithmic approaches.
@@
The Non-Trivial Scaling of Stability: The scaling analysis (Fig. 3) revealed one of the most profound findings of this study: the relationship between network size $N$ and the optimal kernel width $\gamma$ is non-trivial. Contrary to the conventional $\gamma=1/N$ rule of thumb, we found that maintaining optimal storage capacity requires $\gamma$ to be scaled such that the factor $c=\gamma N$ increases with $N$. This indicates that larger networks necessitate more localized kernels to effectively manage the increasingly complex web of inter-pattern interactions in a high-dimensional state space. In essence, as the memory space becomes more crowded with patterns in a higher-dimensional state space, each stored memory must become more “introverted,” narrowing its field of influence to avoid interfering with its neighbors. This scale-dependent requirement for locality is a new, fundamental principle for designing high-capacity kernel-based associative memories. As a practical guideline derived from our findings, optimal performance for random patterns can be achieved by selecting $\gamma$ such that the scaling factor $c=\gamma N$ is in the range of approximately 2 to 5. Once this principle is respected, we confirmed that the storage capacity follows the expected linear scaling law, $P\propto N$, demonstrating the model’s soundness as a scalable architecture.
@@ 6 Conclusion
We used Gemini 2.5 Pro for proofreading the English manuscript.

### v4

@@ Abstract
Kernel-based learning methods such as Kernel Logistic Regression (KLR) can substantially increase the storage capacity of Hopfield networks, but the principles governing their performance and stability remain largely uncharacterized. This paper presents a comprehensive quantitative analysis of the attractor landscape in KLR-trained networks to establish a solid foundation for their design and application. Through extensive, statistically validated simulations, we address critical questions of generality, scalability, and robustness. Our comparative analysis shows that KLR and Kernel Ridge Regression (KRR) exhibit similarly high storage capacities and clean attractor landscapes under typical operating conditions, suggesting that this behavior is a general property of kernel regression methods, although KRR is computationally much faster. We identify a non-trivial, scale-dependent law for the kernel width $\gamma$, demonstrating that optimal capacity requires $\gamma$ to be scaled such that $\gamma N$ increases with network size $N$. This finding implies that larger networks require more localized kernels, in which each pattern’s influence is more spatially confined, to mitigate inter-pattern interference. Under this optimized scaling, we provide clear evidence that storage capacity scales linearly with network size ($P\propto N$). Furthermore, our sensitivity analysis shows that performance is remarkably robust with respect to the choice of the regularization parameter $\lambda$. Collectively, these findings provide a concise set of empirical principles for designing high-capacity and robust associative memories and clarify the mechanisms that enable kernel methods to overcome the classical limitations of Hopfield-type models.
@@ 1 Introduction
Associative memory (AM) is a fundamental cognitive function that enables the retrieval of stored information based on partial or noisy cues. Artificial neural networks, particularly the Hopfield network [1], have long served as foundational models for AM due to their biologically plausible recurrent dynamics and the interpretation of stored memories as stable states (attractors) of an energy function. The recall process in a Hopfield network can be viewed as the network state evolving toward a local minimum of its energy landscape, corresponding to the stored pattern that is most similar to the input cue.
Despite their conceptual elegance and biological grounding, traditional Hopfield networks trained with the Hebbian learning rule [1] exhibit severe limitations. Their storage capacity is theoretically bounded at a pattern-to-neuron ratio $(P/N)$ of approximately 0.14 [2]. Exceeding this limit leads to catastrophic interference, recall errors, and the proliferation of spurious attractors, which do not correspond to any learned pattern yet still act as fixed points of the network dynamics [3]. These spurious attractors compete with desired memories, reduce their basins of attraction, and significantly impair noise robustness.
To overcome these limitations, several approaches have reframed associative memory learning as a supervised optimization problem in which synaptic weights are tuned to make desired patterns stable fixed points. Linear methods, such as the pseudoinverse rule [4] and Linear Logistic Regression (LLR) [5], provide moderate improvements in storage capacity and robustness. However, their performance remains fundamentally constrained by the requirement of linear separability, either in the input space or in the neuron’s input–output mapping.
Nonlinear kernel methods offer a powerful means to transcend these constraints. Techniques such as Support Vector Machines (SVMs), Kernel Ridge Regression (KRR), and Kernel Logistic Regression (KLR) implicitly map data into high- or even infinite-dimensional feature spaces, where complex nonlinear relationships become linearly separable [6]. Applying this framework to associative memory learning holds the potential to dramatically increase storage capacity and enhance noise robustness.
@@ 2 Related Work
While these important lines of research, ranging from classical KAM to MHNs, have demonstrated the potential for high-capacity associative memory, their focus has primarily been on the static properties of the energy function or on performance in specific pattern reconstruction tasks. A detailed and quantitative understanding of the attractor landscape and recall dynamics that arise from practical supervised kernel methods such as KLR and KRR remains largely unexplored. Key questions persist: What happens to spurious attractors in these high-capacity regimes? What are the primary failure modes of recall? What are the geometric properties of the attractors themselves?
The present study aims to address this gap. We connect the high-level capacity analysis of KAM and MHN with the classical dynamical systems perspective on Hopfield networks. By conducting a detailed quantitative analysis of the attractor landscape formed by KLR and comparing it with KRR, for which we have separately demonstrated strong performance and computational advantages [8], we provide a microscopic and dynamical explanation of how these kernel-based methods achieve their notable performance, thereby complementing the existing body of research.
@@ 3.2.1 Hebbian Learning
$s_{i}(t+1)=\text{sign}\left(\sum_{j\neq i}W_{ij}s_{j}(t)\right),$ (2)
where $\text{sign}(z)=1$ if $z\geq 0$ and -1 if $z<0$.
@@ 3.2.4 Kernel Ridge Regression (KRR) Hopfield Network
where $\bm{K}$ is the $P\times P$ kernel (Gram) matrix $K_{\mu\nu}=K(\bm{\xi}^{\mu},\bm{\xi}^{\nu})$, $\bm{I}$ is the identity matrix, and $\bm{Y}$ is the $P\times N$ matrix of stored patterns. This non-iterative solution makes the training process of KRR significantly faster than the gradient-based optimization required for KLR, a difference we quantify in our experiments.
@@ 3.4 Recall Process
The recall process in KLR-trained networks differs from that in Hebbian and LLR-based networks, as it does not rely on an explicit $N\times N$ weight matrix. Given a network state $\bm{s}(t)$ at time $t$, the next state $\bm{s}(t+1)$ is determined by the sign of each neuron’s activation potential relative to a threshold $\bm{\theta}$:
@@ 4.1 Performance and Attractor Landscape Characteristics ($N=500$)
Finally, the recall dynamics are exceptionally fast, as illustrated in Fig. 1(d). For high-similarity initial states, the network consistently converges in only one or two steps. Even for highly corrupted states in the low-similarity regime, the average number of steps to convergence remains remarkably small, typically well below 10 steps. Notably, for these noisy inputs, the convergence time exhibits a non-monotonic “hump” shape, peaking at intermediate storage loads. The subsequent decrease in convergence time at very high loads corresponds precisely to the regime where recall failures dominate (cf. Fig. 1(b)).
This observation suggests that the dynamics rapidly settle into nearby stable states, which correspond to incorrect learned patterns or spurious fixed points, rather than conducting a prolonged search for the correct target. Such rapid convergence represents a meaningful practical advantage of the model.
@@ 4.2 Comparison with Kernel Ridge Regression ($N=100$)
The results of this comparison are summarized in Fig. 2. We first evaluated the recall performance under a moderate noise condition (initial similarity = 0.6), as shown in Fig. 2(a). The performance of the two methods is remarkably similar. Both KLR and KRR maintain near-perfect recall for storage loads up to $P/N\approx 3.0$. Beyond this critical capacity, both methods exhibit a sharp yet smooth degradation in performance, with their recall curves closely aligned. Although minor quantitative differences exist, with KLR displaying a slightly wider shoulder in this case, the overall qualitative behavior is essentially identical. These findings strongly suggest that the observed high-capacity and robust attractor dynamics are not specific to KLR but are likely a more general property of Hopfield-type networks trained with kernel regression methods that effectively separate patterns in a high-dimensional feature space.
@@ 5 Discussion
We hypothesize that two aspects of KLR learning are crucial. First, the kernel trick implicitly maps the input patterns into a high-dimensional feature space where they become more easily separable [6]. This enhanced separability allows the learning algorithm to find decision boundaries that create wide, well-defined basins of attraction for the stored patterns, effectively pushing potential spurious minima away from the recall trajectories. Second, the logistic loss function, combined with L2-regularization, encourages the formation of a “smooth” energy landscape. This property is akin to the margin maximization in SVMs; by pushing the decision boundaries away from the data points, the learning process effectively deepens the basins of attraction around the stored patterns. Unlike Hebbian learning, which is based on simple pairwise correlations and can create many noisy local minima, the KLR objective seeks a solution with a large margin, which tends to flatten the landscape in regions far from the decision boundaries. This landscape smoothing further inhibits the formation of unintended spurious attractors. A detailed theoretical investigation of this mechanism is the subject of our companion paper [12].
@@
Generality of High-Capacity Kernel-Based Memories: Our direct comparison between KLR and KRR (Fig. 2) yields a crucial insight: the remarkable ability to form a clean, high-capacity attractor landscape is not a feature unique to the logistic loss function of KLR. The qualitatively similar performance of KRR suggests that this is a more general property of kernel regression methods applied to Hopfield-type networks. The key mechanism appears to be the kernel trick itself, which maps patterns into a high-dimensional feature space where they become more easily separable, allowing the learning algorithm to sculpt well-defined attractor basins.
Theoretically, this similarity is further supported by the connection between the two algorithms. It is well-established that the optimization of the KLR objective via Newton’s method (Iteratively Reweighted Least Squares) is mathematically equivalent to solving a sequence of weighted Kernel Ridge Regression problems [13, 14, 15]. In this view, KLR is an iteratively refined version of KRR that focuses on difficult patterns. The fact that standard KRR performs comparably suggests that for the random patterns used in this study, such iterative refinement yields only marginal gains. While their performance is similar, the significant computational advantage of KRR’s non-iterative solution makes it a compelling practical alternative, highlighting a clear trade-off between algorithmic approaches.
@@
The Non-Trivial Scaling of Stability: The scaling analysis (Fig. 3) revealed one of the most profound findings of this study: the relationship between network size $N$ and the optimal kernel width $\gamma$ is non-trivial. Contrary to the conventional $\gamma=1/N$ rule of thumb, we found that maintaining optimal storage capacity requires $\gamma$ to be scaled such that the factor $c=\gamma N$ increases with $N$. This indicates that larger networks necessitate more localized kernels to effectively manage the increasingly complex web of inter-pattern interactions in a high-dimensional state space. In essence, as the memory space becomes more crowded with patterns in a higher-dimensional state space, each stored memory must become more “introverted,” narrowing its field of influence to avoid interfering with its neighbors. This scale-dependent requirement for locality is a new, fundamental principle for designing high-capacity kernel-based associative memories. The requirement for locality can also be interpreted as a means to maintain the effective rank of the kernel matrix high enough to ensure linear separability in the high-dimensional feature space. As a practical guideline derived from our findings, optimal performance for random patterns can be achieved by selecting $\gamma$ such that the scaling factor $c=\gamma N$ is in the range of approximately 2 to 5. Once this principle is respected, we confirmed that the storage capacity follows the expected linear scaling law, $P\propto N$, demonstrating the model’s soundness as a scalable architecture.
@@ 6 Conclusion
We used Gemini 2.5 Pro and ChatGPT (GPT-5.1) for proofreading the English manuscript.